Determine the y-coordinate of the centroid of the shaded area. - 13 - - - 13 13 x = kya - - - - - - L - - - 18
Answer #1

te <- î I 72-13 + XX13 consider a defferential Element of width elx and height (Y2-y,) at a distance se from y-axis, - 131 ar718 +[18418 + 1 cup 13x18 + line < 1892) A=195in2 moment of differential Element about we axis dOx = 147+ y 2 ) da 2 = (8,+905 : 3 [1293 + 108 2 + 169x316 69 । 2 / 324 । 169 . (18)2 18 २ 7 3 8- 2788.5 in 3 = 3 - 27885 ♡ -14.3) = y coordinate centor